The angle between the lines 2x = 3y = - z and 6x = - y = - 4z is ?

Grade:12

2 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
Angle between the lines 2x = 3y = - z & 6x = -y = -4z is 90°
Since a1a2 + b1b2 + c1c2 = 0.
where a1=1/2,a2=1/6,b1=1/3,b2=-1,c1=-1,c2=-1/4
Thanks and Regards
Shaik Aasif
askIITians faculty
Rishi Sharma
3 years ago
Dear Student,
Please find below the solution to your problem.

Angle between the lines 2x = 3y = - z & 6x = -y = -4z is 90°
Since a1a2 + b1b2 + c1c2 = 0.
where a1=1/2, a2=1/6,
b1=1/3, b2=-1,
c1=-1, c2=-1/4

Thanks and Regards